Changes

Answer: The five locations on the distant screen have these probabilities:
Line 7: Line 7:  
== Simple case ==
 
== Simple case ==
   −
Consider a double-slit experiment where a particle has a 50% chance of being found, if observed, to be passing through one of the slits.  When it ends up on a distant background screen could then be one of three places for each slit:  further left, straight ahead, or further right, each with a probability of one third.  Assume that the placement on the distant screen of "further right" for one slit is the same as "further left" for the other slit.
+
Consider a double-slit experiment where a particle has a 50% chance of being found, if observed, to be passing through one of the slits.  When it ends up on a distant background screen could then be one of three places for each slit:  further left, straight ahead, or further right, each with a probability of one third.  Assume that the placement on the distant screen of "further right" for one slit is the same as "further left" for the other slit. ''''What is the pattern observed on the distant screen when the double slits remain unobserved?'''''
 +
 
 +
'''''Answer''''':  The five locations on the distant screen have these probabilities:
 +
 
 +
:far left:  1/2 * 1/3 = 1/6
 +
:center left:  1/2 * 1/3 = 1/6
 +
:center:  1/2 * 1/3 * 2 = 1/3
 +
:center right:  1/2 * 1/3 = 1/6
 +
:far right:  1/2 * 1/3 = 1/6
 +
 
 +
:Total = 1
    
==See also==
 
==See also==
Anthony, Irenaeus, Siteadmin, apieditor, Bureaucrats, Check users, nsAm_Govt_101RO, nsAm_Govt_101RW, nsAm_Govt_101_ta, nsJudgesRO, nsJudgesRW, nsJudges_talkRO, nsJudges_talkRW, nsTeam2RO, nsTeam2RW, nsTeam2_talkRO, nsTeam2_talkRW, oversight, Administrators
126,505

edits